OpenStudy (anonymous):

a,b,c,d are 4 consecutive terms of a geometric sequence. Which of the following must be true? I. b^2=ac II. b/a=d/c III. d/a=(c/b)^3 @Loujoelou @phi @Callisto

OpenStudy (phi):

work with the definition of G.P. re write a,b,c,d as a, k a, k^2 a, k^3 a and see what works

OpenStudy (anonymous):

umm..

OpenStudy (phi):

for choice A replace b with k a replace c with k^2 a in b^2 = ac what do you get ?

OpenStudy (anonymous):

ac=k^2 a^2 b=k^2 a^2 the same

OpenStudy (phi):

so A is good. try choice B

OpenStudy (phi):

* choice II

OpenStudy (anonymous):

b/a = ka/a = k d/c = (k^3 a) / (k^2 /a) =k the same

OpenStudy (phi):

is there a typo? I see II. b/d=d/c but you did b/a

OpenStudy (phi):

If II. is b/d = d/c it does not work. if it is b/a = d/c it is true did you try III ?

OpenStudy (anonymous):

d/a = (k^3) /a (c/b)^3 = (k^2a)/(ka) = k they are different, aren't they?

OpenStudy (phi):

you are getting sloppy d = k^3 a so \[ \frac{d}{a} = \frac{k^3 a}{a} = \frac{k^3 \cancel{a}}{\cancel{a}} = k^3\] what do you get for c/b ? then raise that to the 3rd power because you want (c/b)^3

OpenStudy (anonymous):

oh, yes, i miss it. (c/b)^3 = k^3

OpenStudy (phi):

so III. is true

OpenStudy (anonymous):

all are true then

OpenStudy (anonymous):

how about arithmetic sequence? the same way? (different formula)

OpenStudy (phi):

are you saying II. has a typo, and is not correct as given in the question ?

OpenStudy (anonymous):

why? b/a = k d/c = k...

OpenStudy (phi):

yes, but the question asks II. b/d=d/c

OpenStudy (anonymous):

ohh, i typed it wrongly.

OpenStudy (phi):

ok

OpenStudy (phi):

** how about arithmetic sequence? ** none of I. II. or III. will work for A.P.

OpenStudy (anonymous):

nonono, i mean that's another question

OpenStudy (anonymous):

If a,b,c,d are consecutive terms of an arithmetic sequence, which of the following must be true? I. b-a=d-c II. d,c,b,a are consecutive terms of an arithmetic sequence. III. a<b<c<d

OpenStudy (phi):

can you make this a new post. This one is too long.